if A:B = 3:4 and B:C=6:7, then what is C:A?
Answer by amarjeeth123(570)

Given that: A:B = 3:4 and B:C=6:7
(A/B)=(3/4) and (B/C)=(6/7)
(A/C)=(A/B)(B/C)=(3/4)*(6/7)=(9/14)
(C/A)=(14/9)